A steam boiler has a total volume of 3m3.the boiler initially contains 2m3 of saturated steam and 1m3 of saturated liquid at 3mpa.calculaye the mass of vapour, mass of liquid, quality of steam and specific enthalpy of steam

1
Expert's answer
2022-01-03T14:22:53-0500

"m_1=\frac{V_1}{V_1+V_2}\rho pR=0.4 ~kg,"

"m_2=\frac{V_2}{V_1+V_2}\rho pR=0.8 ~kg,"

"m=m_1+m_2=1.2~kg,"

"h=\frac HM=u+pv=7.3~\frac J{mol}."